# Quartet: Native FP4 Training Can Be Optimal for Large Language Models

### v3

@@ 1 Introduction
Over the past decade, the capabilities of large language models (LLMs) have surged, unlocking state‑of‑the‑art performance in AI reasoning, coding, and multimodal understanding. These advances have come at the cost of an unprecedented rise in compute costs, as the floating‑point operations (FLOPs) required to train a frontier model have been doubling every few months [cottier2024rising].
One key lever for reducing compute costs is lower-precision computation: executing the matrix-multiplication (MatMul) kernels that dominate training workloads at lower bit‑widths yields near‑linear gains in throughput and energy efficiency. On the inference side, it is known that 4‑bit quantization—or even lower—can preserve accuracy, via sophisticated calibration and rotation schemes [frantar2022gptq; ashkboos2024quarot; chee2023quip]. For training, recent work has pushed the precision frontier from FP16 [micikevicius2018mixedprecisiontraining] to 8-bit pipelines, responsible in part for efficiency breakthroughs such as DeepSeek‑V3 [deepseekv3]. In this context, NVIDIA’s Blackwell architecture introduces efficient hardware support for even lower-precision microscaling formats [mxfp] such as MXFP and NVFP, which natively support 4‑bit floating‑point operations at higher teraFLOP‑per‑watt efficiency: for instance, moving from 8‑ to 4‑bit multiplies on the B200 GPU can almost double arithmetic throughput, while cutting energy roughly in half [Blackwell].
Yet, today’s algorithmic support for accurate end‑to‑end training in such low precision is missing. State-of-the-art quantized training methods such as Switchback [switchback], Jetfire [jetfire], HALO [ashkboos2025halohadamardassistedlowerprecisionoptimization], and INT4-Transformers [xi2023int4] either (i) lose precision and stability when training current models in 4-bit formats, or (ii) fall back to higher precision for selected matrix multiplications. Bridging this gap calls for both a deeper understanding of quantization error during back‑propagation and new algorithmic safeguards tailored to hardware‑native FP4 formats.
@@
In this paper, we address this challenge via a first systematic study of hardware‑supported FP4 training, focusing on the high-efficiency of the MXFP4 format [mxfp; Blackwell]. Based on this analysis, we introduce an algorithm for MXFP4 native training—in which all matrix multiplications occur in MXFP4—called Quartet, which provides the best accuracy-efficiency trade-off among existing methods, and is near-lossless for LLM pre-training in the large-data regime. Our main technical contribution is a highly-efficient GPU implementation of Quartet, which achieves speedups of almost 2x relative to FP8 for linear layer computations on an NVIDIA Blackwell RTX 5090 GPU. One key achievement is that Quartet enables MXFP4 precision to be “optimal” on the accuracy-efficiency trade-off: at a fixed computational budget, the accuracy impact of lower-precision training in Quartet is fully compensated by the higher efficiency of our implementation. In more detail, our contributions are as follows:
@@
    We apply this framework to MXFP4 precision, seeking to determine if there are practical settings under which native training in this precision can be optimal on Blackwell GPUs. We isolate an algorithm, called Quartet, which achieves this by maximizing both parameter and data efficiency, building on previous SOTA methods for QAT [panferov2025queststabletrainingllms] and quantized backward-pass optimization [tseng2025trainingllmsmxfp4]. Our key technical contribution is a complex, highly-efficient GPU implementation of Quartet specialized to the new Blackwell architecture.
@@
    We validate our approach experimentally by pre-training Llama-family [touvron2023llama2openfoundation] models on the C4 dataset [raffel2020exploring]. Our experiments show that 1) Quartet provides superior accuracy relative to prior methods [xi2023int4; xi2024jetfire; ashkboos2025halohadamardassistedlowerprecisionoptimization] across different computing budgets and model sizes, and that 2) its fast implementation allows it to outperform highly-optimized FP8 kernels. This establishes that MXFP4 can indeed provide “optimal” training in practice.
@@ 2 Related Work
Early work on low-precision neural network training focused on 8-bit or higher precisions, mainly on CNNs. Banner et al. [banner2018scalable] demonstrated accurate 8-bit training via careful scaling and higher-precision accumulation. Yang et al. [yang2020wageubn] proposed a framework that quantized weights, activations, gradients, errors, and even optimizer states to INT, achieving for the first time completely integer-only training with comparable accuracy. SwitchBack [wortsman2023stable] and JetFire [xi2024jetfire] build on this progress, targeting 8-bit training for Transformers [vaswani2017attention]. Specifically, SwitchBack uses a hybrid INT8/BF16 linear layer for vision-language models, performing forward and input-gradient MatMuls in INT8 while computing weight gradients in 16-bit; this yielded 13–25% end-to-end speedups on CLIP models with accuracy within 0.1% of full precision.
JetFire [xi2024jetfire] achieved fully INT8 training for Transformers by using a novel per-block quantization scheme to handle activation and gradient outliers. By partitioning matrices into small blocks and scaling each block independently, JetFire preserved accuracy comparable to FP16 training while obtaining $\sim 40\%$ end-to-end speedup and $1.49\times$ reduction in memory usage. The JetFire approach is conceptually similar to the FP8 DeepSeek training technique [deepseekv3], which used larger block sizes. Recently, HALO [ashkboos2025halohadamardassistedlowerprecisionoptimization] improved upon JetFire in terms of the accuracy-speedup trade-off in INT8, specifically focusing on low-precision fine-tuning. In our work, we will treat FP8 as the idealized baseline that has the quality of BF16 and the speed of raw FP8 GEMM operations. That is, when comparing agains FP8, we compare against simultaneously the most accurate and the fastest FP8-based methods could ever be.
@@
As our results and prior work suggest, going below 8-bit precision in training using the above approaches is extremely challenging, due to the narrower dynamic range and higher error. This frontier was first explored by sun2020ultra, who achieved 4-bit training on ResNets by using a custom numeric format, which unfortunately is far from being supported in hardware. Chmiel et al. [chmiel2023accurate] introduced a logarithmic unbiased quantization (LUQ) scheme to this end, combining two prior ideas: (1) a log-scale FP4-type format to cover a wider dynamic range, and (2) applying stochastic unbiased rounding on the backward. For reference, LUQ incurs a $1.1\%$ top-1 accuracy drop on ResNet50/ImageNet, and has not been validated on hardware-supported FP formats. xi2023int4 proposed a method to train Transformers using INT4 effective precision for all linear layers, using specialized quantizers: block-wise Hadamard transform and LSQ [esser2019learned] for outlier mitigation on the forward pass, and leverage score sampling on the backward pass to exploit structured sparsity, together with a custom INT4-effective format. Their approach trains BERT-family models within 1-2% accuracy gap relative to FP16, with a 2.2x speedup on individual matrix multiplies (relative to 4x theoretical speedup), leading to up to 35% faster training end-to-end.
@@
Given the importance of inference cost reductions, there has been significant work on quantization-aware training (QAT) [choi2018pact; bhalgat2020lsqplus; esser2019learned; baskin2021uniq; wang2023bitnet; kaushal2024spectra], i.e., methods that only quantize the forward pass. Two key difficulties in this setting are 1) minimizing the error induced by quantization on the forward pass, and 2) obtaining a stable gradient estimator over the resulting discrete space. With regards to error reduction, existing methods either try to find a good “learnable” fit w.r.t. the underlying continuous distribution [choi2018pact; esser2019learned], or perform noise injection during QAT in order to make the network more robust to quantization [baskin2021uniq]. Closer to our work, Wang et al. [wang2024fp4] explored FP4 QAT, introducing a “smoother” gradient estimator, together with outlier clamping and compensation to handle activation outliers. While their approach shows good accuracy, it is fairly complex and not validated in terms of efficient support. Prior work by [panferov2025queststabletrainingllms] provided a simpler alternative approach, based on more precise MSE fitting, an optional Hadamard rotation, and a clipping-aware “trust” gradient estimator. By contrast with these forward-only approaches, recent work by tseng2025trainingllmsmxfp4 investigated backward-only quantization with the MXFP4 format, signaling the importance of stochastic rounding and outlier mitigation in low-precision backpropagation.
@@ 3 Background
Quantization maps high-precision internal model states, such as weights, activations, or gradients, to a lower-precision discrete set—i.e., the quantization grid. This grid can be uniform, e.g., for integer quantization, or non-uniform, e.g., floating-point (FP) quantization, where the value spacing is roughly exponential for fixed exponent. Since the original values may differ in scale compared to the grid, a higher-precision scale $s$ is typically stored alongside the quantized values. For a vector $x$, the quantization process can be written as $q(x)=\text{round}\left(\frac{x}{s};\text{grid}\right)$, and the original values can be approximately reconstructed as $\hat{x}=s\cdot q(x)$. Common choices for the scale are setting it to the maximum absolute value (absmax) in $x$ (to avoid clipping) or optimizing it to minimize the mean squared quantization error, e.g. [panferov2025queststabletrainingllms].
@@
Apart from grid choice, quantization methods also differ in the granularity of the scales. A single scale value can be shared across an entire tensor, e.g. [ashkboos2025halohadamardassistedlowerprecisionoptimization], across each row or column [panferov2025queststabletrainingllms], or over more fine-grained custom-defined blocks, such as 22D blocks [jetfire; deepseekv3] or 11D blocks [mxfp; tseng2025trainingllmsmxfp4]. Notably, the latest Blackwell GPU architecture [Blackwell] introduces hardware support for MXFP4/6/8 and NVFP4 formats. MXFP [mxfp] formats share an FP8 power-of-two scale over each 11D block of $32$ elements, while NVFP4 [Blackwell] uses FP8 (E4M3) scales and 1D blocks of $16$ elements.
@@
Quantization typically involves rounding, e.g., via deterministic rounding to the nearest grid point, results in the lowest mean squared error (MSE). In contrast, stochastic rounding introduces randomness, rounding up or down with probabilities based on the input’s distance to nearby grid points. While it may introduce higher MSE, stochastic rounding helps control bias, which can be crucial for maintaining the convergence of iterative optimization algorithms [alistarh2017qsgd].
@@
One key issue when quantizing neural networks is the existence of large outlier values in the network weights, activations, and gradients [dettmers2022llmint8]. One standard way of mitigating such outliers [suresh2017distributed; chee2023quip; ashkboos2025halohadamardassistedlowerprecisionoptimization; ashkboos2024quarot; tseng2025trainingllmsmxfp4] is via the Hadamard transform: given a vector $x\in\mathbb{R}^{d}$, $h(x)$ is defined as $h(x)=H_{d}x$, where $H_{d}\in\mathbb{R}^{d\times d}$ is the normalized Hadamard matrix with elements from $\{\pm 1\}$. Hadamard matrices have a recursive structure $H_{d}=\frac{1}{\sqrt{2}}H_{2}\otimes H_{d/2}$, which enables efficient computation when $d$ is a power of two [fino1976unified]. Optimized FWHT implementations for GPUs are available [dao2024fast; hadacore]. When $d$ is not a power of two, the input vector $x$ is typically either zero-padded to the next power of two or transformed using a Grouped Hadamard Transform, where $x$ is split into equal-sized blocks (each with power-of-two length), and the Hadamard transform is applied independently to each block.
@@
NVIDIA’s 5th-gen. Tensor Cores in Blackwell [Blackwell] provide native 4-bit floating-point execution. The cores support different block-scaled formats such as MXFP4 [mxfp] and NVFP4 [Blackwell], which roughly double the peak throughput over FP8/FP6, with a single B200 GPU peaking at $18$ PFLOPS of dense FP4 compute [Blackwell]. Interestingly, our investigation shows that, as of now, MXFP4 is the only microscaling format with support for all required layouts for both forward and backward multiplications in low precision on Blackwell [Thakkar_CUTLASS_2023]. Therefore, we adopt MXFP4 for our implementation. This format stores each value using 1 sign bit + 1 mantissa bit + 2-bits for exponent. Every group of $32$ elements shares a common 8-bit scaling factor, represented with 8 exponent bits, and no bits for mantissa. Blackwell’s 5th-gen. Tensor Cores handle the required on-the-fly rescaling in hardware, without the need for software-based rescaling at CUDA level. Additional details are provided in Section 4.4.
@@
We pre-train Transformers [vaswani2023attentionneed] of the Llama-2 [touvron2023llama2openfoundation] architecture in the range of 30, 50, 100, 200 million non-embedding parameters across a wide range of data-to-parameter ratios raging from 25x (around compute-optimal [hoffmann2022trainingcomputeoptimallargelanguage]) to 800x (extreme data saturation). We additionally selectively scale the model size up to around 7 billion parameters to verify training stability. We train all models on the train split of the C4 [dodge2021documentinglargewebtextcorpora] dataset and report C4 validation loss as the main metric. We use the AdamW optimizer [loshchilov2019decoupledweightdecayregularization] with weight decay of 0.1, gradient clipping of 1.0, a 10% LR warmup and cosine schedule. We identify the optimal LR for one of the small unquantized baseline models, scale it inverse-proportionally to the number of non-embedding parameters and reuse for every quantization scheme we evaluate. We present all hyper-parameters in Appendix A.1.
@@ 4.1 Ingredient 1: Comparing Quantized Training Approaches via their Induced Scaling Laws
The ability of LLMs to scale predictably with both model size and data across orders of magnitude is a cornerstone of the current AI scaling landscape [kaplan2020scalinglawsneurallanguage]. Mathematically, this says that the expected loss is a function of model and data parameters, often described in the form of a parametric function. This function can be fitted on a set of training runs, and then used to determine the optimal computational training regime [hoffmann2022trainingcomputeoptimallargelanguage] or to extrapolate model performance [grattafiori2024llama3herdmodels].
@@
The key addition is given by the fitted parameters $\text{eff}_{N}(P_{\mathord{\it forward}})$, representing the parameter efficiency of the precision $P_{\mathord{\it forward}}$ used in the forward pass, and $\text{eff}_{D}(P_{\mathord{\it backward}})$ representing the “data efficiency” of the backward pass occurring in a potentially different precision $P_{\mathord{\it backward}}$. (Both these factors are naturally in the interval $(0,1]$, where the value $1$ is reached for full-precision.) Specifically, our parametrization postulates that the impact of the forward-pass precision is felt primarily w.r.t. the trainable parameters, i.e., lowering precision to $P_{\mathord{\it forward}}$ lowers the model’s “effective” parameter count to $N\cdot\text{eff}_{N}(P_{\mathord{\it forward}})\leq N$. This follows the general trend of modeling the effect of forward pass quantization as a multiplicative factor on parameter count [frantar2023scalinglawssparselyconnectedfoundation; kumar2024scalinglawsprecision; frantar2025compressionscalinglawsunifyingsparsity; panferov2025queststabletrainingllms]. For the data term, we postulate that lowering backward-pass precision primarily impacts the data term $D$, so we effectively need additional data to reach the same the same loss, precisely by a factor of $1/\text{eff}_{D}(P_{\mathord{\it backward}})$. This is a novel way to model backward pass quantization that we propose, consistent with optimization theory results [alistarh2017qsgd], as well as observed performance gaps (see Figure 1 (a)). We present experimental data to justify these assumptions and compare against alternative scaling laws [kumar2024scalinglawsprecision] in Appendix A.2.
Experimentally, we observe that different quantized training methods, e.g., STE [STE] vs. QuEST [panferov2025queststabletrainingllms], induce different scaling laws, and in particular different efficiency parameters. While, usually, scaling laws are used to extrapolate model performance across different parameter and data sizes, we propose to use scaling laws to compare different training methods. Specifically, we say that quantized training method A is superior to method B if it offers both higher parameter efficiency $\text{eff}_{N}$ and higher data efficiency $\text{eff}_{D}$.

### v4

@@ 1 Introduction
Over the past decade, the capabilities of large language models (LLMs) have surged, unlocking state‑of‑the‑art performance in AI reasoning, coding, and multimodal understanding. These advances have come at the cost of an unprecedented rise in compute costs, as the floating‑point operations (FLOPs) required to train a frontier model have been doubling every few months [16].
One key lever for reducing compute costs is lower-precision computation: executing the matrix-multiplication (MatMul) kernels that dominate training workloads at lower bit‑widths yields near‑linear gains in throughput and energy efficiency. On the inference side, it is known that 4‑bit quantization—or even lower—can preserve accuracy, via sophisticated calibration and rotation schemes [24; 2; 9]. For training, recent work has pushed the precision frontier from FP16 [33] to 8-bit pipelines, responsible in part for efficiency breakthroughs such as DeepSeek‑V3 [31]. In this context, NVIDIA’s Blackwell architecture introduces efficient hardware support for even lower-precision microscaling formats [36] such as MXFP and NVFP, which natively support 4‑bit floating‑point operations at higher teraFLOP‑per‑watt efficiency: for instance, moving from 8‑ to 4‑bit multiplies on the B200 GPU can almost double arithmetic throughput, while cutting energy roughly in half [35].
Yet, today’s algorithmic support for accurate end‑to‑end training in such low precision is missing. State-of-the-art quantized training methods such as Switchback [54], Jetfire [56], HALO [3], and INT4-Transformers [58] either (i) lose precision and stability when training current models in 4-bit formats, or (ii) fall back to higher precision for selected matrix multiplications. Bridging this gap calls for both a deeper understanding of quantization error during back‑propagation and new algorithmic safeguards tailored to hardware‑native FP4 formats.
@@
In this paper, we address this challenge via a first systematic study of hardware‑supported FP4 training, focusing on the high-efficiency of the MXFP4 format [36; 35]. Based on this analysis, we introduce an algorithm for MXFP4 native training—in which all matrix multiplications occur in MXFP4—called Quartet, which provides the best accuracy-efficiency trade-off among existing methods, and is near-lossless for LLM pre-training in the large-data regime. Our main technical contribution is a highly-efficient GPU implementation of Quartet, which achieves speedups of almost 2x relative to FP8 for linear layer computations on an NVIDIA Blackwell RTX 5090 GPU. One key achievement is that Quartet enables MXFP4 precision to be “optimal” on the accuracy-efficiency trade-off: at a fixed computational budget, the accuracy impact of lower-precision training in Quartet is fully compensated by the higher efficiency of our implementation. In more detail, our contributions are as follows:
@@
    We apply this framework to MXFP4 precision, seeking to determine if there are practical settings under which native training in this precision can be optimal on Blackwell GPUs. We isolate an algorithm, called Quartet, which achieves this by maximizing both parameter and data efficiency, building on previous SOTA methods for QAT [37] and quantized backward-pass optimization [47]. Our key technical contribution is a complex, highly-efficient GPU implementation of Quartet specialized to the new Blackwell architecture.
@@
    We validate our approach experimentally by pre-training Llama-family [46] models on the C4 dataset [39]. Our experiments show that 1) Quartet provides superior accuracy relative to prior methods [58; 57; 3] across different computing budgets and model sizes, and that 2) its fast implementation allows it to outperform highly-optimized FP8 kernels. This establishes that MXFP4 can indeed provide “optimal” training in practice.
@@ 2 Related Work
Early work on low-precision neural network training focused on 8-bit or higher precisions, mainly on CNNs. Banner et al. [4] demonstrated accurate 8-bit training via careful scaling and higher-precision accumulation. Yang et al. [59] proposed a framework that quantized weights, activations, gradients, errors, and even optimizer states to INT, achieving for the first time completely integer-only training with comparable accuracy. SwitchBack [55] and JetFire [57] build on this progress, targeting 8-bit training for Transformers [50]. Specifically, SwitchBack uses a hybrid INT8/BF16 linear layer for vision-language models, performing forward and input-gradient MatMuls in INT8 while computing weight gradients in 16-bit; this yielded 13–25% end-to-end speedups on CLIP models with accuracy within 0.1% of full precision.
JetFire [57] achieved fully INT8 training for Transformers by using a novel per-block quantization scheme to handle activation and gradient outliers. By partitioning matrices into small blocks and scaling each block independently, JetFire preserved accuracy comparable to FP16 training while obtaining $\sim 40\%$ end-to-end speedup and $1.49\times$ reduction in memory usage. The JetFire approach is conceptually similar to the FP8 DeepSeek training technique [31], which used larger block sizes. Recently, HALO [3] improved upon JetFire in terms of the accuracy-speedup trade-off in INT8, specifically focusing on low-precision fine-tuning. In our work, we will treat FP8 as the idealized baseline that has the quality of BF16 and the speed of raw FP8 GEMM operations. That is, when comparing agains FP8, we compare against simultaneously the most accurate and the fastest FP8-based methods could ever be.
@@
As our results and prior work suggest, going below 8-bit precision in training using the above approaches is extremely challenging, due to the narrower dynamic range and higher error. This frontier was first explored by Sun et al. [42], who achieved 4-bit training on ResNets by using a custom numeric format, which unfortunately is far from being supported in hardware. Chmiel et al. [11] introduced a logarithmic unbiased quantization (LUQ) scheme to this end, combining two prior ideas: (1) a log-scale FP4-type format to cover a wider dynamic range, and (2) applying stochastic unbiased rounding on the backward. For reference, LUQ incurs a $1.1\%$ top-1 accuracy drop on ResNet50/ImageNet, and has not been validated on hardware-supported FP formats. Xi et al. [58] proposed a method to train Transformers using INT4 effective precision for all linear layers, using specialized quantizers: block-wise Hadamard transform and LSQ [20] for outlier mitigation on the forward pass, and leverage score sampling on the backward pass to exploit structured sparsity, together with a custom INT4-effective format. Their approach trains BERT-family models within 1-2% accuracy gap relative to FP16, with a 2.2x speedup on individual matrix multiplies (relative to 4x theoretical speedup), leading to up to 35% faster training end-to-end.
@@
Given the importance of inference cost reductions, there has been significant work on quantization-aware training (QAT) [14; 7; 20; 5; 52; 29], i.e., methods that only quantize the forward pass. Two key difficulties in this setting are 1) minimizing the error induced by quantization on the forward pass, and 2) obtaining a stable gradient estimator over the resulting discrete space. With regards to error reduction, existing methods either try to find a good “learnable” fit w.r.t. the underlying continuous distribution [14; 20], or perform noise injection during QAT in order to make the network more robust to quantization [5]. Closer to our work, Wang et al. [53] explored FP4 QAT, introducing a “smoother” gradient estimator, together with outlier clamping and compensation to handle activation outliers. While their approach shows good accuracy, it is fairly complex and not validated in terms of efficient support. Prior work by [37] provided a simpler alternative approach, based on more precise MSE fitting, an optional Hadamard rotation, and a clipping-aware “trust” gradient estimator. By contrast with these forward-only approaches, recent work by Tseng et al. [47] investigated backward-only quantization with the MXFP4 format, signaling the importance of stochastic rounding and outlier mitigation in low-precision backpropagation. Since the first publication of this work, a number of alternative methods for FP4 pre-training have been proposed, either as concurrent alternatives [13], or extensions to Quartet [34; 10].
@@ 3 Background
Quantization maps high-precision internal model states, such as weights, activations, or gradients, to a lower-precision discrete set—i.e., the quantization grid. This grid can be uniform, e.g., for integer quantization, or non-uniform, e.g., floating-point (FP) quantization, where the value spacing is roughly exponential for fixed exponent. Since the original values may differ in scale compared to the grid, a higher-precision scale $s$ is typically stored alongside the quantized values. For a vector $x$, the quantization process can be written as $q(x)=\text{round}\left(\frac{x}{s};\text{grid}\right)$, and the original values can be approximately reconstructed as $\hat{x}=s\cdot q(x)$. Common choices for the scale are setting it to the maximum absolute value (absmax) in $x$ (to avoid clipping) or optimizing it to minimize the mean squared quantization error, e.g. [37].
@@
Apart from grid choice, quantization methods also differ in the granularity of the scales. A single scale value can be shared across an entire tensor, e.g. [3], across each row or column [37], or over more fine-grained custom-defined blocks, such as 22D blocks [56; 31] or 11D blocks [36; 47]. Notably, the latest Blackwell GPU architecture [35] introduces hardware support for MXFP4/6/8 and NVFP4 formats. MXFP [36] formats share an FP8 power-of-two scale over each 11D block of $32$ elements, while NVFP4 [35] uses FP8 (E4M3) scales and 1D blocks of $16$ elements.
@@
Quantization typically involves rounding, e.g., via deterministic rounding to the nearest grid point, results in the lowest mean squared error (MSE). In contrast, stochastic rounding introduces randomness, rounding up or down with probabilities based on the input’s distance to nearby grid points. While it may introduce higher MSE, stochastic rounding helps control bias, which can be crucial for maintaining the convergence of iterative optimization algorithms [1].
@@
One key issue when quantizing neural networks is the existence of large outlier values in the network weights, activations, and gradients [18]. One standard way of mitigating such outliers [43; 9; 3; 2; 47] is via the Hadamard transform: given a vector $x\in\mathbb{R}^{d}$, $h(x)$ is defined as $h(x)=H_{d}x$, where $H_{d}\in\mathbb{R}^{d\times d}$ is the normalized Hadamard matrix with elements from $\{\pm 1\}$. Hadamard matrices have a recursive structure $H_{d}=\frac{1}{\sqrt{2}}H_{2}\otimes H_{d/2}$, which enables efficient computation when $d$ is a power of two [22]. Optimized FWHT implementations for GPUs are available [17; 44]. When $d$ is not a power of two, the input vector $x$ is typically either zero-padded to the next power of two or transformed using a Grouped Hadamard Transform, where $x$ is split into equal-sized blocks (each with power-of-two length), and the Hadamard transform is applied independently to each block.
@@
NVIDIA’s 5th-gen. Tensor Cores in Blackwell [35] provide native 4-bit floating-point execution. The cores support different block-scaled formats such as MXFP4 [36] and NVFP4 [35], which roughly double the peak throughput over FP8/FP6, with a single B200 GPU peaking at $18$ PFLOPS of dense FP4 compute [35]. Interestingly, our investigation shows that, as of now, MXFP4 is the only microscaling format with support for all required layouts for both forward and backward multiplications in low precision on Blackwell [45]. Therefore, we adopt MXFP4 for our implementation. This format stores each value using 1 sign bit + 1 mantissa bit + 2-bits for exponent. Every group of $32$ elements shares a common 8-bit scaling factor, represented with 8 exponent bits, and no bits for mantissa. Blackwell’s 5th-gen. Tensor Cores handle the required on-the-fly rescaling in hardware, without the need for software-based rescaling at CUDA level. Additional details are provided in Section 4.4.
@@
We pre-train Transformers [51] of the Llama-2 [46] architecture in the range of 30, 50, 100, 200 million non-embedding parameters across a wide range of data-to-parameter ratios raging from 25x (around compute-optimal [27]) to 800x (extreme data saturation). We additionally selectively scale the model size up to around 7 billion parameters to verify training stability. We train all models on the train split of the C4 [19] dataset and report C4 validation loss as the main metric. We use the AdamW optimizer [32] with weight decay of 0.1, gradient clipping of 1.0, a 10% LR warmup and cosine schedule. We identify the optimal LR for one of the small unquantized baseline models, scale it inverse-proportionally to the number of non-embedding parameters and reuse for every quantization scheme we evaluate. We present all hyper-parameters in Appendix A.2.
@@ 4.1 Ingredient 1: Comparing Quantized Training Approaches via their Induced Scaling Laws
The ability of LLMs to scale predictably with both model size and data across orders of magnitude is a cornerstone of the current AI scaling landscape [28]. Mathematically, this says that the expected loss is a function of model and data parameters, often described in the form of a parametric function. This function can be fitted on a set of training runs, and then used to determine the optimal computational training regime [27] or to extrapolate model performance [21].
@@
The key addition is given by the fitted parameters $\text{eff}_{N}(P_{\mathord{\it forward}})$, representing the parameter efficiency of the precision $P_{\mathord{\it forward}}$ used in the forward pass, and $\text{eff}_{D}(P_{\mathord{\it backward}})$ representing the “data efficiency” of the backward pass occurring in a potentially different precision $P_{\mathord{\it backward}}$. (Both these factors are naturally in the interval $(0,1]$, where the value $1$ is reached for full-precision.) Specifically, our parametrization postulates that the impact of the forward-pass precision is felt primarily w.r.t. the trainable parameters, i.e., lowering precision to $P_{\mathord{\it forward}}$ lowers the model’s “effective” parameter count to $N\cdot\text{eff}_{N}(P_{\mathord{\it forward}})\leq N$. This follows the general trend of modeling the effect of forward pass quantization as a multiplicative factor on parameter count [26; 30; 25; 37]. For the data term, we postulate that lowering backward-pass precision primarily impacts the data term $D$, so we effectively need additional data to reach the same the same loss, precisely by a factor of $1/\text{eff}_{D}(P_{\mathord{\it backward}})$. This is a novel way to model backward pass quantization that we propose, consistent with optimization theory results [1], as well as observed performance gaps (see Figure 1 (a)). We present experimental data to justify these assumptions and compare against alternative scaling laws [30] in Appendix A.3.
Experimentally, we observe that different quantized training methods, e.g., STE [6] vs. QuEST [37], induce different scaling laws, and in particular different efficiency parameters. While, usually, scaling laws are used to extrapolate model performance across different parameter and data sizes, we propose to use scaling laws to compare different training methods. Specifically, we say that quantized training method A is superior to method B if it offers both higher parameter efficiency $\text{eff}_{N}$ and higher data efficiency $\text{eff}_{D}$.